OpenStudy (anonymous):

find, rank, nullity,row space, column space, null space and left null space of A

OpenStudy (anonymous):

\[\left[ -2, -2, -1, 0 , 1 \right]\] \[\left[ 1,2, 0, 1 , 0 \right]\] \[\left[ -1, 0, -1, 1 , 1 \right]\]

OpenStudy (loser66):

first of all, you have to get rref of A, show me your work

OpenStudy (anonymous):

1 0 1 -1 -1 0 1 -.5 1 .5 0 0 0 0 0

OpenStudy (loser66):

good, how many pivots do you have?

OpenStudy (anonymous):

2

OpenStudy (loser66):

yeap, so? the rank of A is?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

row space is 1st 2 rows

OpenStudy (loser66):

of what, show me those row

OpenStudy (anonymous):

1 0 1 -1 -1 0 1 -.5 1 .5

OpenStudy (loser66):

yep . now, which column do you have pivots?

OpenStudy (loser66):

what are they look like, right on rref

OpenStudy (anonymous):

i dont know

OpenStudy (loser66):

take a look at first 2 column of rref

OpenStudy (loser66):

are they \[\left(\begin{matrix}1 \\ 0\\0\end{matrix}\right) and \left(\begin{matrix}0 \\ 1\\0\end{matrix}\right)\] they are your column space

OpenStudy (loser66):

@ terenze, @ asmistre64 too fun, many people stop by and leave, give me the medal for my stuff???? no help???

OpenStudy (loser66):

@amistre64 @terenzreignz

OpenStudy (amistre64):

medals are up to the discretion of the asker and are not to be "demanded". medals are also given out by bystanders who appreciate the quality of help you provide. My question to medal ratio is about 1:2

OpenStudy (amistre64):

... or is that medal to question ratio?

terenzreignz (terenzreignz):

questions... asked or answered? :D

OpenStudy (loser66):

the Asker runs, just both you and me here..hahaha... too fun,!!

OpenStudy (amistre64):

you did good one the answering part ;) and i consider my questions asked portion to be immaterial ;)

terenzreignz (terenzreignz):

Now you've made me consider my own ratios, @amistre64 :>